OpenStudy (anonymous):

The x-intercept of the graph of ƒ(x)=3log(x+5)-2is: A. 10^(-2/3)+5 B. 10^(-2/3)-5 C. 10^(2/3)+5 D. 10^(2/3)-5

OpenStudy (anonymous):

Set y to 0. You will get 2/3 = log (x+5). Log has base of 10. So we know that for some 10^(2/3) = x+5. 10^(2/3) - 5 is the root if you isolate x!

OpenStudy (anonymous):

thanks!